# Unsupervised cell segmentation by fast Gaussian Processes

### v2

@@ Abstract
Cell boundary information is crucial for analyzing cell behaviors from time-lapse microscopy videos. Existing supervised cell segmentation tools, such as ImageJ, require tuning various parameters and rely on restrictive assumptions about the shape of the objects. While recent supervised segmentation tools based on convolutional neural networks enhance accuracy, they depend on high-quality labelled images, making them unsuitable for segmenting new types of objects not in the database. We developed a novel unsupervised cell segmentation algorithm based on fast Gaussian processes for noisy microscopy images without the need for parameter tuning or restrictive assumptions about the shape of the object. We derived robust thresholding criteria adaptive for heterogeneous images containing distinct brightness at different parts to separate objects from the background, and employed watershed segmentation to distinguish touching cell objects. Both simulated studies and real-data analysis of large microscopy images demonstrate the scalability and accuracy of our approach compared with the alternatives.
@@ 1 Introduction
The spatial organization of cells, such as orientation, density, and shape, are fundamental to characterizing physical mechanisms and biological functions eisenbarth2019dendritic ; prasad2019cell , as the arrangement of cells influences their functions, including proliferation and differentiation liu2007cellular . For example, high cell density can signal cells to die to maintain tissue homeostasis eisenhoffer2012crowding , and cell shape has been linked to DNA synthesis and cell growth folkman1978role . Advances in microscopy techniques enable detailed video recordings of cellular motions. Image segmentation approaches have been developed for microscopy images meijering2012cell , where the segmentation results, including cell locomotion, alignment, and density, can be used for modeling cell behaviors selmeczi2005cell ; gu2023data ; fang2024inverse ; khang2025automated . Therefore, precise cell image segmentation is essential for quantifying cell location, alignment, and morphology, as these statistics enhance our understanding of how cellular spatial organization influences human health outcomes tsukui2020collagen ; barbazan2023cancer .
@@
Cellular segmentation tools can be broadly classified as unsupervised and supervised methods, based on whether labeled data is needed for training the models. One of the most popular tools of unsupervised cellular segmentation is the ImageJ abramoff2004image ; schneider2012nih ; schindelin2012fiji , which enables segmenting both cell nuclei and whole cells, and linking cells across frames of a microscopy video via plug-ins such as TrackMate tinevez2017trackmate . Prior to using ImageJ for cell detection and segmentation, a series of user-driven image pre-processing steps must be completed in the program. For instance, the image is typically converted into a binary image first, based on the threshold chosen either manually or automatically via several global threshold setting methods nichele2020quantitative ; liu2009otsu . Additional image processing steps, such as despeckling for noise reduction ferreira2011imagej , typically need to be run before segmentation. To detect cell objects within an image, the algorithm scans the corresponding binary image until the edge of an object is reached. Then, the algorithm traces around the edge of the object until the starting point is reached. This process is repeated for pixels of each object, or the foreground pixels, in the binary image. The centroid of each cell object can be identified by fitting an ellipse for each object using the second-order central moments schneider2012nih , which assumes the cell boundary is elliptical. The parameters for the best-fit ellipse, including the major and minor axes and centroids, and matrices for the object masks for each object boundary, are recorded in ImageJ ferreira2011imagej . Additional functionality, such as extension of elliptical outlines, can be added to the user interface by writing macro code to streamline image analysis or adding third-party plug-ins schneider2012nih . However, achieving reasonable segmentation results is often labor-intensive and requires human interaction at each step. Furthermore, the assumption of the shape of the cells in ImageJ is restrictive. Though the ellipse can approximate nuclei of certain cells, it cannot approximate the shape of cytoplasm or the whole cell of frequently used cells in experiments, such as fibroblasts or epithelial cells.
@@
Supervised cellular segmentation methods, are typically built upon a convolutional neural network structure, such as the U-net ronneberger2015u . A popular tool is $\sf{Cellpose}$, which builds U-nets for segmenting images that contain one type of cells that are trained on a dataset of over 70,000 segmented objects, including annotated images that contain numerous types of cells stringer2021cellpose . The model first compresses the image and then reconstructs the image to its original dimensions through a U-net to capture both local and global features. It further generates a topological map from diffusion simulation to product gradient maps for the model to learn the horizontal and vertical gradients of the image. The final layer of Cellpose produces three feature maps, including the horizontal and vertical directional gradients, and the sigmoid transformation of the probability map of the output image from the U-net, and the parameters of the Cellpose are then estimated by fitting a loss function of the gradient vector and binary label of each pixel of the image. As the Cellpose segmentation requires massive training data, it may not work well for detecting novel cell types, where high-labeled image data that may not be readily available. Furthermore, training Cellpose and using it for segmenting large images with several million pixels can be computationally expensive.
@@
In response to the challenges the current methods face, we present a novel unsupervised cell image segmentation method for segmenting images containing one type of cells, based on separable Gaussian processes, automated thresholding, and watershed operations for image segmentation. The proposed method has several advantages. First, the separable covariance structure on image data enables us to decompose the large covariance into a Kronecker product of two small covariances, which substantially reduces the computational cost for computing the matrix inversion and the logarithm of the determinant of the covariance matrix. Second, unlike conventional unsupervised image segmentation methods such as ImageJ, which require manual tuning of multiple hyperparameters, including Gaussian blur radius, thresholding method, and minimum object size, our proposed segmentation pipeline is free of manual parameter selection. We developed fast algorithms for maximum likelihood estimation of the parameters in GPs and the predictive distribution. The predictive mean surface naturally facilitates foreground-background separation by applying a data-driven threshold with the optimal threshold automatically estimated based on the absolute second difference in foreground pixel counts between the thresholds. Third, the cell masks are segmented using the watershed algorithm, which simulates water flowing through a topographical map of the image to separate cell objects. Fourth, our approach is entirely unsupervised, negating any need for annotated training data, in contrast to the supervised methods, such as Cellpose. Thus, the proposed method is particularly suitable for segmenting new cell types or other biological structures. We studied our methodology by simulated data for noise filtering. We also compared our approaches with the alternatives using real microscopy images for segmenting both the cell nucleus and whole cell in optically dense images, a challenging scenario even when labeled data are available. These studies demonstrate that our approach substantially improves the alternatives.
@@
The rest of the article is organized as follows. We discuss GP models of images and a fast algorithm for parameter estimation and predictive distribution in Section 2. In Section 3, we introduce our workflow for cellular image segmentation from Gaussian processes, including image smoothing, binary thresholding, and segmentation. In Section 4, we used simulated data and microscopy images of cells to compare our methods and other alternatives for denoising, including principal component analysis and dynamic mode decomposition. We also demonstrate the computational advantage of the fast algorithm against direct computation. Finally, we generate cell masks for both the cell nucleus and cytoplasm, or the whole cell, by our methods and ImageJ segmentation method to compare their accuracy with the manual labels in Section 5. Our findings not only present a novel cell segmentation technique but also highlight the applicability of fast Gaussian processes in image processing tasks. The data and code of this article are made publicly available: https://github.com/UncertaintyQuantification/cell_segmentation.
@@ 2.1 The likelihood function and predictive distribution
Let us consider a Gaussian process (GP) model for a two-dimensional (2D) microscopy image, where the intensity at the pixel location $\mathbf{x}$ is modeled as $y(\mathbf{x})=f(\mathbf{x})+\epsilon$, with $f(\cdot)$ being a latent Gaussian process (GP) and $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{0})$ being a Gaussian noise with variance $\sigma^{2}_{0}$ rasmussen2006gaussian . Consider an $n_{1}\times n_{2}$ image $\mathbf{Y}$, where the $(i,j)$th term is $y(\mathbf{x}_{i,j})$ for the pixel location $\mathbf{x}_{i,j}=(x_{i,1},x_{j,2})$. The data matrix can be written as:
@@
where $\mathbf{F}$ follows a matrix normal distribution $\mathbf{F}\sim\mbox{Matrix-Normal}\left(\mu\mathbf{1}_{n_{1}\times n_{2}},\,\sigma^{2}\mathbf{R}_{1},\mathbf{R}_{2}\right)$ with $\mu$ being a mean parameter, $\mathbf{1}_{n_{1}\times n_{2}}$ being a $n_{1}\times n_{2}$ matrix of ones, $\sigma^{2}$ being the variance parameter, $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ being correlation matrices for two inputs, respectively, and the vectorized noise matrix follows $\mbox{Vec}(\mathbf{E})\sim\mathcal{MN}(0,\sigma^{2}_{0}\mathbf{I}_{N\times N})$, with $N=n_{1}n_{2}$. Additional trend structure can be modeled in the mean function, and the covariance can be generalized to be semi-separable in the model gu2022gaussian .
@@
where $\mathbf{\tilde{R}}=\mathbf{R}+\eta\mathbf{I}_{N}$, $\mathbf{R}=\mathbf{R}_{2}\otimes\mathbf{R}_{1}$, $\otimes$ is a Kronecker product and $\eta=\sigma^{2}_{0}/\sigma^{2}$ is a nugget parameter. Here the correlation matrices are parameterized kernel functions $R_{1}(i,i^{\prime})=K_{1}(x_{i,1},x_{i^{\prime},1})$ and $R_{2}(j,j^{\prime})=K_{2}(x_{j,2},x_{j^{\prime},2})$. Frequently used kernel functions include the power exponential kernel and Matérn kernel rasmussen2006gaussian . Denote $K_{l}(x_{i,l},x_{i^{\prime},l})=K_{l}(d_{l})$ with $d_{l}=|x_{i,l}-x_{i^{\prime},l}|$. For instance, the Matérn kernel with roughness parameter $5/2$ follows handcock1993bayesian :
@@
where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable rasmussen2006gaussian , and it is used as a default choice of some software packages of GP surrogate models roustant2012dicekriging ; gu2018robustgasp . We use the kernel function in (3) for demonstration purposes.
After specifying the model, we compute the likelihood function and predictive distribution for parameter estimation and prediction, respectively. The parameters of the model in (2) include the mean, variance, range and nugget parameters $\{\mu,\sigma^{2},\bm{\gamma},\eta\}$ with $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ being the range parameters in the kernel function in Equation (3). As the number of pixels is large, we employed the maximum likelihood estimator (MLE) to estimate the parameters, as the result is similar to the robust estimation by marginal posterior mode Gu2018robustness . Given the range and nugget parameters, the MLE of the mean and variance parameters has a closed-form expression: $\hat{\mu}=(\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N})^{-1}\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$, where $\hat{\sigma}^{2}=\frac{S^{2}}{N}$ with $S^{2}=(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})^{T}\bm{\tilde{}}{\mathbf{R}}^{-1}(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$. The logarithm of the profile likelihood after plugging the MLE of the mean and variance parameters follows:
@@
Vectorize signal $\mathbf{f}_{v}=\mbox{vec}(\mathbf{F})$ from model (1). After plugging the MLE of the parameters, the posterior distribution of the image vector follows follows
@@ 2.2 Fast computation for images
A wide range of approximation approaches of GPs have been developed snelson2006sparse ; lindgren2011explicit ; rue2009approximate ; datta2016hierarchical ; katzfuss2017multi ; gramacy2015local ; guinness2017circulant ; kaufman2008covariance ; zhu2024radial . For GPs with product covariances on image data, no approximation is required. Denote the eigendecomposition of the subcovariance $\mathbf{R}_{l}=\mathbf{U}_{l}\bm{\Lambda}_{l}\mathbf{U}^{T}_{l}$, where $\mathbf{U}_{l}$ and $\bm{\Lambda}_{l}$ are matrices of eigenvectors and a diagonal matrix of eigenvalues for $l=1,2$, respectively. Furthermore, denote $\bm{\lambda}_{l}=(\lambda_{1,l},...,\lambda_{n_{l},l})^{T}$, an $n_{l}$-vector of the diagonal terms in $\bm{\Lambda}_{l}$, for $l=1,2$.
@@ 3 Image segmentation from Gaussian processes
Figure 1: Workflow for segmenting and labeling cell images: (A) Divide the image into different sub-images to enable locally estimated mean and variance parameters for capturing local properties such as change of brightness. (B) Compute the predictive mean of fast GPs in Section 2.2 to each sub-image, which greatly reduces image noise. (C) Threshold each smoothed sub-image based on the criterion discussed in Section 3.2 to produce binary images, separating cells from the background. The optimal threshold is estimated for each sub-image. (D) Recombine the binary sub-images into a single binary image, and apply the watershed algorithm discussed in Section 3.3 to the image for segmentation and labeling, with each cell marked by a unique color.
Though Gaussian processes have been widely used in as surrogate models to emulate expensive computer simulations, nonparametric regression, and modeling spatially correlated data rasmussen2006gaussian ; chang2016calibrating , the flexibility and scalability of GPs in unsupervised image segmentation have not been studied yet. The overall flow of our cellular image segmentation process is summarized in Figure 1. First, some microscopy images of cells are particularly large with $10^{6}-10^{7}$ pixels. Similar to other image segmentation methods schneider2012nih ; stringer2021cellpose , we segment each large image into multiple sub-images. We then apply the fast algorithm of GPs in Lemma 1 to get the predictive mean of each sub-image, illustrated in Parts (A) and (B) in Figure 1, respectively. We assume the range parameters and nugget of the GPs are the same for each subimage estimated by MLE in Lemma 1, yet the mean and variance parameters are distinct for each separable image, and the estimation of these parameters has a closed-form expression. Such a step enables different estimated parameters to capture the change of optical properties, such as brightness difference and out-of-focus blur from distinct parts of a large microscopy image, which are crucial in practice. Second, we developed a robust measure that will be introduced in Section 3.2 to threshold the predictive mean of each sub-image into background and objects, showed in Panel (C). Finally, the binary sub-images are restitched in Panel (D), and the watershed algorithm is applied to the binary matrix to retrieve the labeled cell masks, which will be introduced in Section 3.3.

### v3

@@ Abstract
Cell boundary information is crucial for analyzing cell behaviors from time-lapse microscopy videos. Existing supervised cell segmentation tools, such as ImageJ, require tuning various parameters and rely on restrictive assumptions about the shape of the objects. While recent supervised segmentation tools based on convolutional neural networks enhance accuracy, they depend on high-quality labeled images, making them unsuitable for segmenting new types of objects not in the database. We developed a novel unsupervised cell segmentation algorithm based on fast Gaussian processes for noisy microscopy images without the need for parameter tuning or restrictive assumptions about the shape of the object. We derived robust thresholding criteria adaptive for heterogeneous images containing distinct brightness at different parts to separate objects from the background, and employed watershed segmentation to distinguish touching cell objects. Both simulated studies and real-data analysis of large microscopy images demonstrate the scalability and accuracy of our approach compared with the alternatives.
@@ 1 Introduction
The spatial organization of cells, such as orientation, density, and shape, are fundamental to characterizing physical mechanisms and biological functions Eisenbarth (2019); Prasad and Alizadeh (2019), as the arrangement of cells influences their functions, including proliferation and differentiation Liu and Chen (2007). For example, high cell density can signal cells to die to maintain tissue homeostasis Eisenhoffer et al. (2012), and cell shape has been linked to DNA synthesis and cell growth Folkman and Moscona (1978). Advances in microscopy techniques enable detailed video recordings of cellular motions. Image segmentation approaches have been developed for microscopy images Meijering (2012), where the segmentation results, including cell locomotion, alignment, and density, can be used for modeling cell behaviors Selmeczi et al. (2005); Gu et al. (2023); Fang and Gu (2024); Khang et al. (2025). Therefore, precise cell image segmentation is essential for quantifying cell location, alignment, and morphology, as these statistics enhance our understanding of how cellular spatial organization influences human health outcomes Tsukui et al. (2020); Barbazan et al. (2023).
@@
Cellular segmentation tools can be broadly classified as unsupervised and supervised methods, based on whether labeled data is needed for training the models. One of the most popular tools of unsupervised cellular segmentation is the ImageJ Abràmoff et al. (2004); Schneider et al. (2012); Schindelin et al. (2012), which enables segmenting both cell nuclei and whole cells, and linking cells across frames of a microscopy video via plug-ins such as TrackMate Tinevez et al. (2017). Prior to using ImageJ for cell detection and segmentation, a series of user-driven image pre-processing steps must be completed in the program. For instance, the image is typically converted into a binary image first, based on the threshold chosen either manually or automatically via several global threshold setting methods Nichele et al. (2020); Liu and Yu (2009). Additional image processing steps, such as despeckling for noise reduction Ferreira and Rasband (2011), typically need to be run before segmentation. To detect cell objects within an image, the algorithm scans the corresponding binary image until the edge of an object is reached. Then, the algorithm traces around the edge of the object until the starting point is reached. This process is repeated for pixels of each object, or the foreground pixels, in the binary image. The centroid of each cell object can be identified by fitting an ellipse for each object using the second-order central moments Schneider et al. (2012), which assumes the cell boundary is elliptical. The parameters for the best-fit ellipse, including the major and minor axes and centroids, and matrices for the object masks for each object boundary, are recorded in ImageJ Ferreira and Rasband (2011). Additional functionality, such as extension of elliptical outlines, can be added to the user interface by writing macro code to streamline image analysis or adding third-party plug-ins Schneider et al. (2012). However, achieving reasonable segmentation results is often labor-intensive and requires human interaction at each step. Furthermore, the assumption of the shape of the cells in ImageJ is restrictive. Though the ellipse can approximate nuclei of certain cells, it cannot approximate the shape of cytoplasm or the whole cell of frequently used cells in experiments, such as fibroblasts or epithelial cells.
@@
Supervised cellular segmentation methods, are typically built upon a convolutional neural network structure, such as the U-net Ronneberger et al. (2015). A popular tool is $\sf{Cellpose}$, which builds U-nets for segmenting images that contain one type of cells that are trained on a dataset of over 70,000 segmented objects, including annotated images that contain numerous types of cells Stringer et al. (2021). The model first compresses the image and then reconstructs the image to its original dimensions through a U-net to capture both local and global features. It further generates a topological map from diffusion simulation to product gradient maps for the model to learn the horizontal and vertical gradients of the image. The final layer of Cellpose produces three feature maps, including the horizontal and vertical directional gradients, and the sigmoid transformation of the probability map of the output image from the U-net, and the parameters of the Cellpose are then estimated by fitting a loss function of the gradient vector and binary label of each pixel of the image. As the Cellpose segmentation requires massive training data, it may not work well for detecting novel cell types, where high-labeled image data that may not be readily available. Furthermore, training Cellpose and using it for segmenting large images with several million pixels can be computationally expensive.
@@
In response to the challenges the current methods face, we present a novel unsupervised cell image segmentation method for segmenting images containing one type of cells, based on separable Gaussian processes, automated thresholding, and watershed operations for image segmentation. The proposed method has several advantages. First, the separable covariance structure on image data enables us to decompose the large covariance into a Kronecker product of two small covariances, which substantially reduces the computational cost for computing the matrix inversion and the logarithm of the determinant of the covariance matrix. Second, unlike conventional unsupervised image segmentation methods such as ImageJ, which require manual tuning of multiple hyperparameters, including Gaussian blur radius, thresholding method, and minimum object size, our proposed segmentation pipeline is free of manual parameter selection. We developed fast algorithms for maximum likelihood estimation of the parameters in GPs and the predictive distribution. The predictive mean surface naturally facilitates foreground-background separation by applying a data-driven threshold with the optimal threshold automatically estimated based on the absolute second difference in foreground pixel counts between the thresholds. Third, the cell masks are segmented using the watershed algorithm, which simulates water flowing through a topographical map of the image to separate cell objects. Fourth, our approach is entirely unsupervised, negating any need for annotated training data, in contrast to the supervised methods, such as Cellpose. Thus, the proposed method is particularly suitable for segmenting new cell types or other biological structures. We studied our methodology by simulated data for noise filtering. We also compared our approaches with the alternatives using real microscopy images for segmenting both the cell nucleus and whole cell in optically dense images, a challenging scenario even when labeled data are available. These studies demonstrate that our approach substantially outperforms the alternatives.
@@
The rest of the article is organized as follows. We discuss GP models of images and a fast algorithm for parameter estimation and predictive distribution in Section 2. In Section 3, we introduce our workflow for cellular image segmentation from Gaussian processes, including image smoothing, binary thresholding, and segmentation. In Section 4, we use simulated data and microscopy images of cells to compare our methods and other alternatives for denoising, including principal component analysis and dynamic mode decomposition. We also demonstrate the computational advantage of the fast algorithm against direct computation. Finally, we generate cell masks for both the cell nucleus and cytoplasm, or the whole cell, by our methods and ImageJ segmentation method to compare their accuracy with the manual labels in Section 5. Our findings not only present a novel cell segmentation technique but also highlight the applicability of fast Gaussian processes in image processing tasks. The data and code of this article are made publicly available: https://github.com/UncertaintyQuantification/cell_segmentation.
@@ 2.1 The likelihood function and predictive distribution
Let us consider a Gaussian process (GP) model for a two-dimensional (2D) microscopy image, where the intensity at the pixel location $\mathbf{x}$ is modeled as $y(\mathbf{x})=f(\mathbf{x})+\epsilon$, with $f(\cdot)$ being a latent Gaussian process (GP) and $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{0})$ being a Gaussian noise with variance $\sigma^{2}_{0}$ Rasmussen (2006). Consider an $n_{1}\times n_{2}$ image $\mathbf{Y}$, where the $(i,j)$th term is $y(\mathbf{x}_{i,j})$ for the pixel location $\mathbf{x}_{i,j}=(x_{i,1},x_{j,2})$. The data matrix can be written as:
@@
where $\mathbf{F}$ follows a matrix normal distribution $\mathbf{F}\sim\mbox{Matrix-Normal}\left(\mu\mathbf{1}_{n_{1}\times n_{2}},\,\sigma^{2}\mathbf{R}_{1},\mathbf{R}_{2}\right)$ with $\mu$ being a mean parameter, $\mathbf{1}_{n_{1}\times n_{2}}$ being a $n_{1}\times n_{2}$ matrix of ones, $\sigma^{2}$ being the variance parameter, $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ being correlation matrices for two inputs, respectively, and the vectorized noise matrix follows $\mbox{Vec}(\mathbf{E})\sim\mathcal{MN}(0,\sigma^{2}_{0}\mathbf{I}_{N\times N})$, with $N=n_{1}n_{2}$. Additional trend structure can be modeled in the mean function, and the covariance can be generalized to be semi-separable in the model Gu and Li (2022).
@@
where $\mathbf{\tilde{R}}=\mathbf{R}+\eta\mathbf{I}_{N}$, $\mathbf{R}=\mathbf{R}_{2}\otimes\mathbf{R}_{1}$, $\otimes$ is a Kronecker product and $\eta=\sigma^{2}_{0}/\sigma^{2}$ is a nugget parameter. Here the correlation matrices are parameterized kernel functions $R_{1}(i,i^{\prime})=K_{1}(x_{i,1},x_{i^{\prime},1})$ and $R_{2}(j,j^{\prime})=K_{2}(x_{j,2},x_{j^{\prime},2})$. Frequently used kernel functions include the power exponential kernel and Matérn kernel Rasmussen (2006). Denote $K_{l}(x_{i,l},x_{i^{\prime},l})=K_{l}(d_{l})$ with $d_{l}=|x_{i,l}-x_{i^{\prime},l}|$. For instance, the Matérn kernel with roughness parameter $5/2$ follows Handcock and Stein (1993):
@@
where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable Rasmussen (2006), and it is used as a default choice of some software packages of GP surrogate models Roustant et al. (2012); Gu et al. (2019). We use the kernel function in (3) for demonstration purposes.
After specifying the model, we compute the likelihood function and predictive distribution for parameter estimation and prediction, respectively. The parameters of the model in (2) include the mean, variance, range and nugget parameters $\{\mu,\sigma^{2},\bm{\gamma},\eta\}$ with $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ being the range parameters in the kernel function in Equation (3). As the number of pixels is large, we employed the maximum likelihood estimator (MLE) to estimate the parameters, as the result is similar to the robust estimation by marginal posterior mode Gu et al. (2018). Given the range and nugget parameters, the MLE of the mean and variance parameters has a closed-form expression: $\hat{\mu}=(\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N})^{-1}\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$, where $\hat{\sigma}^{2}=\frac{S^{2}}{N}$ with $S^{2}=(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})^{T}\bm{\tilde{}}{\mathbf{R}}^{-1}(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$. The logarithm of the profile likelihood after plugging the MLE of the mean and variance parameters follows:
@@
Vectorize signal $\mathbf{f}_{v}=\mbox{Vec}(\mathbf{F})$ from model (1). After plugging the MLE of the parameters, the posterior distribution of the image vector follows
@@ 2.2 Fast computation for images
A wide range of approximation approaches of GPs have been developed Snelson and Ghahramani (2006); Lindgren et al. (2011); Rue et al. (2009); Datta et al. (2016); Katzfuss (2017); Gramacy and Apley (2015); Guinness and Fuentes (2017); Kaufman et al. (2008); Zhu et al. (2024). For GPs with product covariances on image data, no approximation is required. Denote the eigendecomposition of the subcovariance $\mathbf{R}_{l}=\mathbf{U}_{l}\bm{\Lambda}_{l}\mathbf{U}^{T}_{l}$, where $\mathbf{U}_{l}$ and $\bm{\Lambda}_{l}$ are matrices of eigenvectors and a diagonal matrix of eigenvalues for $l=1,2$, respectively. Furthermore, denote $\bm{\lambda}_{l}=(\lambda_{1,l},...,\lambda_{n_{l},l})^{T}$, an $n_{l}$-vector of the diagonal terms in $\bm{\Lambda}_{l}$, for $l=1,2$.
@@ 3 Image segmentation from Gaussian processes
Figure 1: Workflow for segmenting and labeling cell images: (A) Divide the image into different sub-images to enable locally estimated mean and variance parameters for capturing local properties such as the change of brightness. (B) Compute the predictive mean of fast GPs in Section 2.2 to each sub-image, which greatly reduces image noise. (C) Threshold each smoothed sub-image based on the criterion discussed in Section 3.2 to produce binary images, separating cells from the background. The optimal threshold is estimated for each sub-image. (D) Recombine the binary sub-images into a single binary image, and apply the watershed algorithm discussed in Section 3.3 to the image for segmentation and labeling, with each cell marked by a unique color.
Though Gaussian processes have been widely used in as surrogate models to emulate expensive computer simulations, nonparametric regression, and modeling spatially correlated data Rasmussen (2006); Chang et al. (2016),the flexibility and scalability of GPs in unsupervised image segmentation have not been studied yet. The overall flow of our cellular image segmentation process is summarized in Figure 1. First, some microscopy images of cells are particularly large with $10^{6}-10^{7}$ pixels. Similar to other image segmentation methods Schneider et al. (2012); Stringer et al. (2021), we segment each large image into multiple sub-images. We then apply the fast algorithm of GPs in Lemma 1 to get the predictive mean of each sub-image, illustrated in Parts (A) and (B) in Figure 1, respectively. We assume the range parameters and nugget of the GPs are the same for each subimage estimated by MLE in Lemma 1, yet the mean and variance parameters are distinct for each separable image, and the estimation of these parameters has a closed-form expression. Such a step enables different estimated parameters to capture the change of optical properties, such as brightness difference and out-of-focus blur from distinct parts of a large microscopy image, which are crucial in practice. Second, we developed a robust measure that will be introduced in Section 3.2 to threshold the predictive mean of each sub-image into background and objects, as shown in Panel (C). Finally, the binary sub-images are restitched in Panel (D), and the watershed algorithm is applied to the binary matrix to retrieve the labeled cell masks, which will be introduced in Section 3.3.